OpenStudy (avai70178):

Please help me. Jason considered two similar televisions at a local electronics store. The generic version was based off the brand name and was three fourths the size of the brand name. If the generic television set is 12 inches by 24 inches, what are the dimensions of the brand name television? 18 inches by 36 inches 6 inches by 12 inches 16 inches by 32 inches 9 inches by 18 inches

OpenStudy (anonymous):

9 inches by 18 inches

OpenStudy (avai70178):

thank yiu!

OpenStudy (anonymous):

you welcome

OpenStudy (whpalmer4):

(3/4)*12=9 (3/4)*24 =18

OpenStudy (anonymous):

correct whpalmer